Agenda/Minutes on 2020/11/20
CRY+VIS Interface Meeting
Participants: Takahashi, Hirata, Yano, Sato, Ikeda, Aso, Ushiba, Tomaru, Akutsu, Matteo
- Evacuation of EYC was started on 18th. Large reak was found. It whould be necessary to open the chamber again.
- The wire was disconnected from the cryo-payload in EX on 19th.
- Schedule: Evacuation of IYC will be started in the morning on 21st.
- Jig for the FC painting: Asking to make 2D drawings. After that, we will go to purchase order.
PF actuator: Measurement of coefficients --> calculation of the noise (by mid Dec.).
Participants: Takahashi, Hirata, Yano, Sato, Fabian, Matteo, Sugiyama, Aso, Terrence, Tanaka, Washimi
- Closed the feedthrough flanges on the top chamber in EY. The chamber was evacuated from 18th.
- Assembled the side stoppers and replaced the side plates on the BF in EY.
- Disconnected the wire from the cryo-payload in EX. The cable for the TF actuator was broken and one magnet dropped into the corrugate tube during this work.
- Released the jack for the top chamber in IY. The top chamber went down by 0.1-0.5mm.
[Type-A paper] Then what is going on about the paper reviewing... My JGW account looks unavailable now and I cannot collect much information about the current publication rules or so. How should I work on ... (Y. Fujii)
[Hirata-san, Washmi-san and Satou-san, Ikeda-san, Fabian]
- They assembled BS F0 yaw stepper motor and SR limit switches.
- They Removed screws from SR3 chamber top and side doors.
- They did cable inventory.
- Ikeda-san finished the script to actuate on the jammed stepper motor.
We checked IP limit switches using the medm screen and physically pushing them with an Allen key. They worked well although they are very stiff. One of them experienced a plastic deformation and had to be bent back. Satou-san system seems to be better. Klog 15484.
- After Yokozawa-san's actions the temperature became 23 degrees, which is suitable.
On Friday I tried to moving the jammed stepper motor when applying a large actuation on the keystone, both ways. It didn't solve the problem. Klog 15479.
- After working this week inside of SRM chamber, we have the impression we may be able to add weight to the SR3 BF without taking it out of the chamber and, at the same time, do it in a suitable way to keep the IRM alignment.
- To add mass to the IM we would require removing only two IRM panels and one OSEM.
- Next week we can inspect the BF and find whether it's possible or not.
SRM trouble mechanical troubleshooting
On Monday I recovered the alignment before in situ inspection in the mine. Klog 15474.
General visual inspection: everything looked fine, including cables. Klog 15474.
Although no cable seemed to be causing problems, for the sake of trying we unclamped from the Flying Saucer the only real suspect cable, which is the picomotor kapton cable. We measured a transfer function and the problem was still there. Klog 15474.
- We tried all the suitable holes at the bottom and at the top (keystone lock screws and nuts access).
- Once inside the BF, we cannot point the camera where we want. Mostly we only saw the cap and sometimes the keystone.
- We were able to identify some components but it's still too dark. We need more light.
- We inspected visually the bottom central section of the BF, which is the coil-magnet actuator assembly. Nothing seemed wrong.
The three of us (one at the time) pulled the suspension rod down by hand, and it moved by several milimeters smoothly. It didn't feel like rubbing anything. Klog 15485.
Assessment of SRM from the control room: see this presentation.
Netx two weeks plan
- According to the Microsoft Project:
- Next week is for limit switches in SR2 and
- The week after the next is for "adjustment of ballast mass in SR3", which is separate a separate item from "Repair removing suspension". They are actually the same.
- For the next two weeks we can:
- Install the limit switches in all suspensions (F0 yaw and vertical).
- Look for accessible places where we could add mass on SR3 BF from within the chamber and write a procedure.
- Fabian still has to design some simple ballast masses for SR3 IM.
- Aso-san can take a fresh look to SRM payload if he wants.
- (No progress)
- Started assembly of the circuits for the combiner. Some parts were lack.
- Total 24ch combiner circuits for Type-A can be procured using the ICRR budget (2.5M JPY) in this FY.
- Type-B remedy work was started.
- Procedure documents.
- Criteria in the acceptance test.
- JPS Meeting 2021 Spring (3/12-15).
The week from 11/23
- Aso: 11/23-27
- Takahashi: 11/24-27
- Hirata: 11/24-26
- Sato: 11/24-27
The week from 11/30
- Takahashi: 11/30-12/4
- Hirata: 11/30-12/4
- On 2020/11/27(Fri)